OpenStudy (anonymous):

Paul walks 8 miles in 2 hours. Find the unit rate

jimthompson5910 (jim_thompson5910):

8 miles/2 hours = (8/2) miles/1 hr = 4 miles/1 hour

jimthompson5910 (jim_thompson5910):

So his unit rate is 4 miles/1 hour which is commonly expressed as 4 miles per hour or 4 mph

OpenStudy (anonymous):

8/2 = x/1 (cross multiply) 2x = 8 x = 4 The unit rate is 4 mph.